OpenStudy (anonymous):

The length of a rectangle is 5cm more than its width. If its perimeter is 58cm,how do you calculate its dimensions and its area?

OpenStudy (anonymous):

width: x length: x+5 perimeter: 2( x+ x + 5) = 58 find width and length then its area

OpenStudy (anonymous):

nphuongsun93 where did you get the 2 from in the calculation of the perimeter?

OpenStudy (anonymous):

the perimeter equation is 2 * (width + length) because there are 2 widths and lengths

OpenStudy (anonymous):

L = W + 5 2L + 2W = 58 L + W = 29 (W + 5) + W = 29 2W = 24 Find W and then L from here...

OpenStudy (anonymous):

ok lets see if im on the right track so far

OpenStudy (anonymous):

i got 12 as the width and 17 as the length

OpenStudy (anonymous):

so 12+12+17+17 =58

OpenStudy (anonymous):

yep sounds good

OpenStudy (anonymous):

Yes, you have done it correctly...

OpenStudy (anonymous):

Ok A= l*w A= 17 *12 so: A = 204

OpenStudy (anonymous):

You have become Star now...

OpenStudy (anonymous):

yay :) thank you!